# Semantic Web: Past, Present, and Future

### v1

@@ 10 Impact for Practitioners
Linking and using graph data on the Web has become a widespread practice. Today, there is a large amount of open data in various formats and domains, such as bibliographic information management, bioinformatics, and e-government. DBpedia is the central hub in this context, around which different datasets and domains are grouped (cf. [24]). This is illustrated, e. g., by the tremendous growth of the Linked Open Data Cloud [^69] The growth of the Linked Open Data Cloud is documented at: http://linkeddata.org/. since 2007. Two of the latest notable supporters of graph-based data are online auctioneer eBay with their graph database [^70] https://github.com/eBay/akutan and the U.S. space agency NASA with the unification of internal distributed case databases as knowledge graphs [^71] https://blog.nuclino.com/why-nasa-converted-its-lessons-learned-database-into-a-knowledge-graph. These and other success stories of the Semantic Web in industries and industry-scale knowledge graphs are described by Noy et al. [118]. Further analyses and surveys arguing about the importance but also challenges of using graph data can be found in the literature like the 2020 survey of Sahu et al. [138] and the 2021 reflection about the future of graphs by Sakr et al. [139]. The usefulness of knowledge graphs and semantic-based data modeling for complex systems is also discussed in the 2024 book by Abonyi et al. [2]. The importance of graph databases is also reflected by the Forbes business magazine, which predicted in 2019 that graph databases would be the next mainstream database technology [^72] https://www.forbes.com/sites/cognitiveworld/2019/07/18/graph-databases-go-mainstream.
@@
Regarding more heavyweight ontologies in OWL, there has also been movement in recent years. In addition to numerous research-derived inference engines such as Pellet and Hermit mentioned above, inference mechanisms for OWL can now be found in commercial graph databases such as neo4j [^82] https://neo4j.com/blog/neo4j-rdf-graph-database-reasoning-engine/. Furthermore, pattern-based core ontologies can also be found in software development workflows [144]. The development and use of core ontologies is part of a continuous delivery process that is used in practice.
@@ 11 Summary and Outlook
  • Neuro-symbolic systems: As mentioned in the introduction, we see as an important direction of future work the combination of symbolic AI and subsymbolic AI. By combining the strength of Large Language Models (LLM), i. e., generative AI, in processing and generating natural language text and accessing structured data and logical reasoning capabilities of the Semantic Web, a next step towards the vision of automated agents that perform complex planning tasks may be reached. An example is performing A* search with an LLM [174]. Specifically, LLMs might comprehensively capture and acquire human knowledge [45], but current LLMs lack responding to simple questions of non-existing facts in their training data [45], may not contain all facts [152], and thus return less accurate answers [91]. To leverage the distinct capabilities of both LLMs and the Semantic Web, the integration of neuro-symbolic systems appears to offer a viable solution [124]. Neuro-symbolic systems could also address the problem that LLMs’ output is based on the most probable answer, which sometimes leads to wrong answers – often referred to as “hallucinations” [17, 88, 152].
@@
Finally, we like to point to existing literature discussing the future directions of research on the Semantic Web [33] and Knowledge Graphs [48]. Breit et al. [28] conducted a survey on the fusion of Semantic Web and Machine Learning, exploring the opportunities arising from the convergence of these two paradigms.

### v2

@@ 12 Impact for Practitioners
Linking and using graph data on the Web has become a widespread practice. Today, there is a large amount of open data in various formats and domains, such as bibliographic information management, bioinformatics, and e-government. DBpedia is the central hub in this context, around which different datasets and domains are grouped (cf. [DBLP:journals/expert/Bizer09]). This is illustrated, e. g., by the tremendous growth of the Linked Open Data Cloud [^68] The growth of the Linked Open Data Cloud is documented at: http://linkeddata.org/. since 2007. Two of the latest notable supporters of graph-based data are online auctioneer eBay with their graph database [^69] https://github.com/eBay/akutan and the U.S. space agency NASA with the unification of internal distributed case databases as knowledge graphs [^70] https://blog.nuclino.com/why-nasa-converted-its-lessons-learned-database-into-a-knowledge-graph. These and other success stories of the Semantic Web in industries and industry-scale knowledge graphs are described by Noy et al. [DBLP:journals/cacm/NoyGJNPT19]. Further analyses and surveys arguing about the importance but also challenges of using graph data can be found in the literature like the 2020 survey of Sahu et al. [DBLP:journals/vldb/SahuMSLO20] and the 2021 reflection about the future of graphs by Sakr et al. [DBLP:journals/cacm/SakrBVIAAAABBDV21]. The usefulness of knowledge graphs and semantic-based data modeling for complex systems is also discussed in the 2024 book by Abonyi et al. [AbonyiEtAl2024OntologiesInIndustry]. The importance of graph databases is also reflected by the Forbes business magazine, which predicted in 2019 that graph databases would be the next mainstream database technology [^71] https://www.forbes.com/sites/cognitiveworld/2019/07/18/graph-databases-go-mainstream.
@@
Regarding more heavyweight ontologies in OWL, there has also been movement in recent years. In addition to numerous research-derived inference engines such as Pellet and Hermit mentioned above, inference mechanisms for OWL can now be found in commercial graph databases such as neo4j [^81] https://neo4j.com/blog/neo4j-rdf-graph-database-reasoning-engine/. Furthermore, pattern-based core ontologies can also be found in software development workflows [DBLP:conf/semweb/SchonteichKS18]. The development and use of core ontologies is part of a continuous delivery process that is used in practice.
@@ 13 Summary and Outlook
  • Neuro-symbolic systems: As mentioned in the introduction, we see as an important direction of future work the combination of symbolic AI and subsymbolic AI. By combining the strength of Large Language Models (LLM), i. e., generative AI, in processing and generating natural language text and accessing structured data and logical reasoning capabilities of the Semantic Web, a next step towards the vision of automated agents that perform complex planning tasks may be reached. An example is performing A* search with an LLM [DBLP:journals/corr/abs-2310-13227-toolchainstar]. Specifically, LLMs might comprehensively capture and acquire human knowledge [10.1145/3608966], but current LLMs lack responding to simple questions of non-existing facts in their training data [10.1145/3608966], may not contain all facts [DBLP:journals/corr/abs-2308-10168], and thus return less accurate answers [Hou2023-fo]. To leverage the distinct capabilities of both LLMs and the Semantic Web, the integration of neuro-symbolic systems appears to offer a viable solution [DBLP:journals/corr/abs-2306-08302]. Neuro-symbolic systems could also address the problem that LLMs’ output is based on the most probable answer, which sometimes leads to wrong answers – often referred to as “hallucinations” [DBLP:journals/corr/abs-2302-04023, DBLP:conf/adbis/Hose23, DBLP:journals/corr/abs-2308-10168].
@@
Finally, we like to point to existing literature discussing the future directions of research on the Semantic Web [DBLP:journals/corr/abs-2309-13939] and Knowledge Graphs [DBLP:journals/corr/abs-2308-14217]. Breit et al. [10.1145/3586163] conducted a survey on the fusion of Semantic Web and Machine Learning, exploring the opportunities arising from the convergence of these two paradigms.